# Consensus-based construction of high-dimensional free energy surface

### v3

@@
Consensus-based construction of high-dimensional free energy surface
Liyao Lyu
Department of Computational Mathematics, Science & Engineering, Michigan State University, MI 48824, USA &Huan Lei
Department of Computational Mathematics, Science & Engineering, Michigan State University, MI 48824, USA
Department of Statistics & Probability, Michigan State University, MI 48824, USA
One essential problem in quantifying the collective behaviors of molecular systems lies in the accurate construction of free energy surfaces (FESs). The main challenges arise from the prevalence of energy barriers and the high dimensionality. Existing approaches are often based on sophisticated enhanced sampling methods to establish efficient exploration of the full phase space. On the other hand, the collection of optimal sample points for the numerical approximation of FESs remains largely under-explored, where the discretization error could become dominant for systems with a large number of collective variables (CVs). We propose a consensus sampling based approach by reformulating the construction as a minimax problem which simultaneously optimizes the function representation and the training set. In particular, the maximization step establishes a stochastic interacting particle system to achieve the adaptive sampling of the max-residue regime by modulating the exploitation of the Laplace approximation of the current loss function and the exploration of the uncharted phase space; the minimization step updates the FES approximation with the new training set. By iteratively solving the minimax problem, the present method essentially achieves an adversarial learning of the FESs with unified tasks for both phase space exploration and posterior error enhanced sampling. We demonstrate the method by constructing the FESs of molecular systems with a number of CVs up to $30$.
Keywords Free energy |||| Coarse-graining |||| Machine learning |||| Molecular simulation |||| Adversarial learning
Molecular dynamics (MD) provides an essential tool [1] to access the micro-scale insights of complex processes in many scientific applications, including chemical engineering, material synthesis, and drug design. To further predict the collective behaviors, it is often desirable to construct the free energy surfaces (FESs) with respect to a set of functional-relevant collective variables (CVs). However, the accurate construction of FES remains a practical challenge due to the high dimensionality and the prevalence of energy barriers. Several ingenious methods based on importance sampling have been developed, such as umbrella sampling [2], histogram reweighting [3], metadynamics [4, 5], variational enhanced sampling [6, 7, 8], weighted ensemble and its variants [9, 10, 11], hyperdynamics [12], orthogonal space random walk [13] and adaptive biasing force  [14, 15, 16]. Despite the efficacy in enhancing the phase space exploration, these methods are based on iterative estimation of the biased probability density functions (PDFs) (or biased force), where the efficiency generally diminishes as the number of CVs increases. Strategies such as bias-exchange [17] and parallel bias metadynamics [18, 19] have been developed to alleviate this issue. Both methods factorize the biasing potential as a product of several low-dimensional functions and can work effectively when the underlying FESs admit such structures which, however, may not always be the case for high-dimensional problems. Several other methods [20, 21, 22, 23, 24, 25, 26, 27, 28] achieve enhanced sampling irrespective of pre-defined CVs (see review [29]). However, the construction of explicit high-dimensional FESs remains a challenging problem.
Alternative to the histogram methods, the temperature-accelerated molecular dynamics [30] and the adiabatic free energy dynamics [31, 32] are developed by introducing an extended dynamics of the CVs with an artificially high temperature to overcome the energy barriers. The methods avoid dealing with the biased potentials and have shown the promise of efficient exploration of the high-dimensional CV space [33]. On the other hand, to further construct the FES, the sampling points are deposited in a greedy way [34]; the effectiveness for high-dimensional problems remains under-explored.
With the recent advancements in machine learning, kernel methods and deep neural networks (DNNs) have demonstrated promising results in representing multi-dimensional FES [35, 36, 37, 8, 38]. However, as also noted in Ref. [39], one essential problem for constructing high-dimensional FES lies in how to optimize the training set. Ideally, to efficiently approximate a high-dimensional function, the sampling should enable certain adaptivity based on the posterior error so that the training set can be updated according to the particular underlying FES. However, such posterior error-based sampling remains largely open for FES construction since the approximation residual can neither be easily queried nor sampled within the full phase space. In practice, the training samples are often collected either as pre-defined grid points or in a heuristic manner. As the number of CVs increases, such choices may induce pronounced approximation error and suffer from performance degradation, since most phase space regimes essentially become thermodynamically inaccessible. Remarkably, the DeepVes [8] enables the efficient sampling of the CV space by jointly constructing the bias potential and the target distribution. On the other hand, the accurate reconstruction of explicit FES further requires the estimation of high-dimensional PDFs from the samples. Recent efforts [40, 41] seek the direct approximation of the transition operator instead of estimating PDFs using generative models, when representative configurations are known a priori. In addition, the reinforced dynamics [42, 43] (see also Ref. [44]) proposes an efficient approach to impose adaptivity by using an uncertainty indicator as an indirect measure of the construction error to bias MD simulations, which relies on calculating the standard deviation of the predictions from multiple DNNs trained on the same dataset.
@@
In this work, we present a consensus-based enhanced sampling (CES) method to efficiently construct high-dimensional FES for complex MD systems. A unique feature is that the method enables adaptive posterior error-based sampling such that the FES approximation and the sample points can be simultaneously optimized. Also, unlike the reinforced dynamics, the method does not rely on the training of multiple DNN representations. The main idea is to formulate the construction as a min-max problem, where the max-problem seeks a residue-based distribution to establish adaptive sampling in the vicinity of the explored phase space regime, and the min-problem optimizes the DNN parameters for the FES representation. Iteratively optimizing both the training set and the DNN representation achieves an adversarial construction of the FES pertaining to the kinetic transition regime. For the maximization step, it is worth mentioning that the query of the thermodynamically accessible phase space is non-trivial. In particular, existing approaches [45, 46] in the spirit of adversarial generative models can not be directly applied, since the global residual error is unknown a prior. Instead, we establish a consensus-based sampling  [47] (see also Ref.  [48, 49, 50, 51] ) in the form of a stochastic interacting particle system governed by a McKean stochastic differential equation. A quadratic potential is adaptively constructed to probe the local maximum error regime by exploiting the Laplace approximation under a low-temperature limit. Meanwhile, a coherent noise term is introduced to efficiently explore the full CV space under a high-temperature limit and yield the target sampling points for the maximization problem, by which the minimization of the DNN residue can be solved jointly. We demonstrate the effectiveness of the proposed method by constructing the FES of several biomolecule systems with the number of CVs up to 30. Figure 1 sketches the workflow of the proposed method.
@@
Figure 1: The workflow of the present CES-based method. In the minimization step, given a collection of sampling points, the reference force (i.e., the gradient of the underlying FES) can be calculated using the restrained dynamics; a comparison with the force inferred from the DNN approximation yields a posterior error. In the maximization step, the posterior error determines a residue-based distribution with entropy regularization. An interacting particle system following McKean stochastic dynamics is used to achieve an adaptive sampling of the max-residue regime by modulating the exploitation of the Laplace approximation of the current residue-based distribution and the exploration of the uncharted phase space. The FES can be accurately reconstructed after several iterations of the minimization and maximization step.
Consensus-based construction method
Free energy and mean forces
@@
$A(\mathbf{z})=-\frac{1}{\beta}\ln\rho(\mathbf{z}),$ (1)
@@
where $\beta=1/k_{B}T$ is the inverse of the thermal temperature, $\rho(\mathbf{z})=\frac{1}{Z}\int\exp{(-\beta U(\mathbf{r}))}\delta(\mathbf{s}(\mathbf{r})-\mathbf{z})\mathrm{d}\mathbf{r}$ is the marginal density for $\mathbf{s}(\mathbf{r})=\mathbf{z}$, $Z=\int\exp{(-\beta U(\mathbf{r}))}\mathrm{d}\mathbf{r}$ is the partition function, and $\delta(\cdot)$ represents the Dirac delta function. For high-dimensional CVs, the direct estimation of $\rho(\mathbf{z})$ often becomes numerically challenging. An alternative approach is to fit the mean force $\mathbf{F}(\mathbf{z}):=-\nabla A(\mathbf{z})$ at various sample points, which can be estimated via the restrained dynamics [52] by introducing a harmonic term into the full potential, i.e.,
$U_{k}(\mathbf{r},\mathbf{z})=U(\mathbf{r})+\frac{k}{2}(\mathbf{s}(\mathbf{r})-\mathbf{z})^{T}(\mathbf{s}(\mathbf{r})-\mathbf{z}),$ (2)
@@
where $k$ represents the magnitude of the restrained potential. As shown in Ref [30], the mean force can be computed by $\mathbf{F}(\mathbf{z})=\lim_{k\to\infty}\mathbf{F}^{k}(\mathbf{z})$, where $\mathbf{F}^{k}(\mathbf{z})$ is defined by
$\mathbf{F}^{k}(\mathbf{z})=\frac{1}{Z_{k}(\mathbf{z})}\int k(\mathbf{s}(\mathbf{r})-\mathbf{z})\exp{(-\beta U_{k}(\mathbf{r},\mathbf{z}))}\mathrm{d}\mathbf{r},$ (3)
and can be sampled as the local first-moment estimation.
@@
In principle, given a collection of sample points $\mathbf{z}$, $A(\mathbf{z})$ can be re-constructed (up to a constant) by matching the mean force $-\nabla A(\mathbf{z})$ at the individual points. However, for complex MD systems, the sampling over the phase space could be highly non-trivial due to the prevalence of local energy minima; the training set is often determined a prior as pre-selected points or in a greedy manner. As the number of CVs increases, the empirical random samples may introduce pronounced discretization error of the loss function over the full phase space. To efficiently construct $A(\mathbf{z})$ in the thermodynamically accessible regime, it is desirable to simultaneously optimize the training set and the FES approximation through certain adaptive sampling based on the posterior residual error. This motivates the present method illustrated below.
Min-Max formulation
@@
Let $A_{\mathcal{N}}(\mathbf{z})$ denote the DNN representation of the FES $A(\mathbf{z})$, which is parameterized by minimizing the loss function
$\mathcal{L}_{\mathcal{N}}(\mathbf{z})=\|\nabla_{\mathbf{z}}A_{\mathcal{N}}(\mathbf{z})+\mathbf{F}(\mathbf{z}))\|^{2}$ (4)
@@
for $\mathbf{z}\in\Gamma$. To solve the problem, we introduce a sampling distribution $q(\mathbf{z})\in W$, where $W$ is a function space that defines a proper constraint of the PDF $q(\mathbf{z})$ on $\Gamma$, and define the weighted loss
$(\mathcal{L}_{\mathcal{N}},q)=\int_{\Gamma}\mathcal{L}_{\mathcal{N}}(\mathbf{z})q(\mathbf{z})\mathrm{d}\mathbf{z}.$ (5)
@@
A desired distribution intends to maximize the discrepancy in the dataset for a given network $A_{\mathcal{N}}(\mathbf{z})$. Accordingly, we define the maximum problem as
$G[\mathcal{L}_{\mathcal{N}}]=\max_{q}(\mathcal{L}_{\mathcal{N}},q).$ (6)
@@
$\min_{A_{\mathcal{N}}}\max_{q}(\mathcal{L}_{\mathcal{N}},q).$ (7)
Proposition 1.
@@
Assuming that there exists a solution of $\mathcal{L}_{\mathcal{N}}(\mathbf{z})=0$ for $\mathbf{z}\in\Gamma$, then $A^{\ast}$ is a solution if and only if it solves (7).
Proof.
@@
Suppose $A^{\ast}$ is the solution for $\mathcal{L}_{\mathcal{N}}(\mathbf{z})=0$, it satisfies $(\mathcal{L}_{\mathcal{N}^{\ast}},q)=0$ for any $q\in W$, i.e., $G[\mathcal{L}_{\mathcal{N}^{\ast}}]=0$. Therefore, $A^{\ast}$ is a solution for the minimax problem (7). On the other hand, if $\hat{A}$ is the minimizer for problem (7) but not the solution for $\mathcal{L}_{\mathcal{N}}(\mathbf{z})=0$, then there exists $\hat{q}\in V$ such that $(\mathcal{L}_{\hat{\mathcal{N}}},\hat{q})>0$. However, $(\mathcal{L}_{\mathcal{N}^{\ast}},q)=0$ for all $q$, which contradicts the assumption that $\hat{A}$ is the minimizer. ∎
@@
Proposition 1 shows that the direct construction of FES $A(\mathbf{z})$ can be reformulated as an adversarial learning of the optimal solution $A_{\mathcal{N}}(\mathbf{z})$ for the min-max problem (7). Accordingly, the training consists of two components: the minimization part optimizes the DNN representation with the current training set; the maximization part explores the regime of the largest residue for the current DNN representation and essentially establishes an adaptive sampling of the training set based on the posterior error.
@@
To numerically solve the max-problem, certain regularization needs to be introduced. Otherwise, the max-problem will simply yield a delta measure, i.e., $\delta(\mathbf{z}-\mathbf{z}^{*})$, where $\mathbf{z}^{\ast}=\arg\max\mathcal{L}_{\mathcal{N}}(\mathbf{z})$. Since the sampling needs to simultaneously identify the max-residual regime and explore the uncharted phase space, we introduce the entropy-based regularization [53, 54] (see also Refs. [55, 56, 45] for gradient-based regularization), and the max-problem is reformulated by
$\min_{q}\int(-\mathcal{L}_{\mathcal{N}}(\mathbf{z})+\kappa_{h}^{-1}\ln q(\mathbf{z}))q(\mathbf{z})\mathrm{d}\mathbf{z}.$ (8)
@@
The problem is convex for a PDF $q$ with a unique global minimum at $q^{*}(\mathbf{z})=\exp(-\kappa_{h}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))/Z^{*}$, where $Z^{*}=\int\exp(-\kappa_{h}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))\mathrm{d}\mathbf{z}$ and $\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z})=-\mathcal{L}_{\mathcal{N}}(\mathbf{z})$. The parameter $\kappa_{h}$ is a Lagrangian multiplier, balancing the focus between the peak concentration and the scope of exploration, and is somewhat similar to the inverse temperature in statistical mechanics. An elevated $\kappa_{h}^{-1}$ induces a distribution closer to a uniform distribution. Conversely, a diminished $\kappa_{h}^{-1}$ induces a concentrated distribution near the max-residue point.
Given $q^{\ast}(\mathbf{z})$, the direct sampling remains challenging for high-dimensional CVs. Inspired by the consensus-based sampling method [47], we establish a stochastic particle system. By properly constructing the conservative potential and noise terms, the particle distribution provides a good approximation of $q^{\ast}$ as illustrated below.
Exploitation and exploration in the max-problem

### v4

@@
Abstract
We present a consensus-based framework that unifies phase space exploration with posterior-residual-based adaptive sampling for surrogate construction in high-dimensional energy landscapes. Unlike standard approximation tasks where sampling points can be freely queried, systems with complex energy landscapes such as molecular dynamics (MD) do not have direct access to arbitrary sampling regions due to the physical constraints and energy barriers; the surrogate construction further relies on the dynamical exploration of phase space, posing a significant numerical challenge. We formulate the problem as a minimax optimization that jointly adapts both the surrogate approximation and residual-enhanced sampling. The construction of free energy surfaces (FESs) for high-dimensional collective variables (CVs) of MD systems is used as a motivating example to illustrate the essential idea. Specifically, the maximization step establishes a stochastic interacting particle system to impose adaptive sampling through both exploitation of a Laplace approximation of the max-residual region and exploration of uncharted phase space via temperature control. The minimization step updates the FES surrogate with the new sample set. Numerical results demonstrate the effectiveness of the present approach for biomolecular systems with up to 30 CVs. While we focus on the FES construction, the developed framework is general for efficient surrogate construction for complex systems with high-dimensional energy landscapes.
keywords:
@@ Abstract
Consensus-based sampling, Adaptive sampling, Minimax optimization, Phase space exploration, High-dimensional free energy
{MSCcodes}
65C35, 65K10, 82C31
1 Introduction
Many problems in computational science rely on the efficient approximation of physical quantities in systems governed by high-dimensional energy landscapes, where direct sampling and evaluation are constrained by complex system dynamics. Canonical examples include the computation of transition paths in gradient systems [49, 13], the construction of coarse-grained molecular dynamics (MD) models [35], and uncertainty quantification under energy-induced probability measures [29]. Unlike the standard surrogate construction tasks where the sampling points can be freely queried, these problems further rely on efficient exploration and sampling over the phase space, where the thermodynamically accessible regions are often unknown a priori. There are two essential challenges: (\@slowromancapi@) the prevalence of energy barriers, which makes direct sampling inefficient and prone to getting trapped in local minima; various enhanced sampling strategies are often required; (\@slowromancapii@) the high dimensionality of the surrogate model, which generally requires a large number of samples and motivates adaptive sampling strategies based on the approximation error of the target quantity. In practice, the efficient surrogate construction should account for both enhanced sampling with the complex energy landscapes and the residual-based adaptivity in constrained phase space — yet simultaneously addressing both remains a nontrivial and open computational challenge.
In this study, we aim to develop a unified approach that enables both efficient phase space exploration and residual-based adaptive sampling for constrained phase space learning problems. As a motivating application, we consider the construction of free energy surfaces (FESs) with respect to a set of collective variables (CVs) for MD systems. While this is a long-standing problem in computational science, the accurate construction of high-dimensional FESs remains a difficult task due to the aforementioned two challenges. Most existing methods primarily target the first challenge based on various importance sampling strategies to overcome energy barriers, such as umbrella sampling [44], histogram reweighting [26], metadynamics [27, 18], variationally enhanced sampling [46, 40, 4], and adaptive biasing force  [10, 11, 28, 9]. Alternatively, temperature-accelerated [32] and adiabatic dynamics [39, 1] introduce an extended dynamics of the CVs with an artificially high temperature to facilitate the phase space exploration. Despite their broad applications, the computational efficiency of these methods generally degrades as the number of CVs increases. More importantly, these methods do not explicitly address the second challenge. The residual error is not incorporated into the adaptive sampling and FES construction, which limits their effectiveness in high-dimensional problems.
From a different perspective, several approaches related to the second challenge have been developed based on adaptive sampling [42, 51, 43, 16, 14, 23, 20] and adversarial learning [52, 3, 53] for solving high-dimensional partial differential equations (PDEs). The essential idea is to introduce certain residual-based distributions or weak formulations, where new collocation points or test basis functions can be adaptively updated during the training process. While they have shown promising results for high-dimensional PDEs, these methods rely on the free query of new sample points and their residual error within the domain. As such, they cannot be directly applied to the present problem, where the global residual error is unknown a priori. In particular, the new sample points cannot be freely placed within the phase space but need to be navigated through dynamical exploration of the thermodynamically accessible regions. Alternatively, the reinforced dynamics (RiD) [54] (see also Ref. [47]) proposes using the uncertainty indicator as a proxy for the residual error to bias MD simulations, which, however, relies on calculating the standard deviation of the predictions from an ensemble of neural network (NN) surrogates trained on the same dataset. Moreover, the phase space exploration is constrained by the underlying energy landscape, which typically requires small time steps due to the stiffness and roughness of the MD potential function.
@@ 1 Introduction
To address the above two challenges, we present a consensus-based adaptive sampling (CAS) method to efficiently construct surrogates within high-dimensional energy landscapes with applications to FES construction in complex MD models. A unique feature is that the present method enables gradient-free residual-based adaptive sampling such that the FES approximation and the phase-space exploration can be simultaneously optimized. The method is formulated as a minimax optimization problem. The max-problem seeks a residual-based distribution to establish adaptive sampling in the vicinity of the explored phase space regime, while the min-problem optimizes the FES surrogate based on the new samples. For the maximization step, we emphasize that the establishment of the residual-based distribution is only formal; the analytical form of the distribution is unknown and the value at an arbitrary point can not be directly obtained. As a result, common sampling approaches based on Markovian Chain Monte Carlo (MCMC) [37] and Langevin dynamics [38] are not applicable, as explained in more detail in Sec. 2.2. Instead, we establish a consensus-based sampling [7] (see also Refs. [6, 8]) in the form of a stochastic interacting particle system governed by a McKean stochastic differential equation (SDE). The gradient-free nature enables us to collect new samples adaptive to the local residual without the analytical form of the target distribution. Specifically, a quadratic potential is adaptively constructed to probe the local max-residual regime by exploiting the Laplace approximation under a low-temperature limit. Meanwhile, a coherent noise term is introduced to efficiently explore the full CV space under a high-temperature limit and yield the updated sampling points used for the subsequent minimization step.
The present iterative procedure achieves adversarial learning of the FES pertaining to the thermodynamically important regions. In contrast to existing adaptive sampling methods for PDE solvers [42, 43, 16], the present method does not rely on the free query of arbitrary sample points. Instead, it enables us to establish a dynamical exploration of the phase space along with the surrogate construction. Moreover, unlike the reinforced dynamics [54] that navigates the sampling points through the biased MD simulations, the present sampling dynamics is governed by a smooth quadratic potential irrespective of the roughness of the underlying energy landscape, enabling much larger time steps and more efficient phase-space exploration. We demonstrate the effectiveness of the proposed method by constructing the FES of biomolecule systems involving up to $30$ CVs. Fig. 1 sketches the workflow of the proposed method.
@@
Figure 1: The workflow of the present CAS-based method. In the minimization step, given a collection of sampling points, the reference force $F(\mathbf{z})$ (i.e., the gradient of the underlying FES) can be calculated using the restrained molecular dynamics; a comparison with the force inferred from the NN surrogate $A_{\mathcal{N}}(\mathbf{z})$ yields the loss function $L_{\mathcal{N}}(\mathbf{z})=|\nabla_{\mathbf{z}}A_{\mathcal{N}}(\mathbf{z})+F(\mathbf{z})|^{2}$. In the maximization step, the loss function determines a residual-based distribution with entropy regularization $q^{\ast}(\mathbf{z})\propto\exp(\kappa_{h}\mathcal{L_{N}}(\mathbf{z}))$. An interacting particle system following the McKean SDE is used to achieve adaptive sampling of the max-residual region by modulating the exploitation of the Laplace approximation of $q^{\ast}(\mathbf{z})$ and the exploration of the uncharted phase space. The FES can be accurately reconstructed after several iterations of the minimization and maximization step.
2 Methods
2.1 Free energy and mean forces
@@
(1) $A(\mathbf{z})=-\frac{1}{\beta}\ln\rho(\mathbf{z}),$
where $\beta=1/k_{B}T$ is the inverse of the thermal temperature,
(2) $\rho(\mathbf{z})=\frac{1}{Z}\int\exp{(-\beta U(\mathbf{r}))}\delta(\mathbf{s}(\mathbf{r})-\mathbf{z})\mathrm{d}\mathbf{r}$
@@
is the marginal probability density function (PDF) for $\mathbf{s}(\mathbf{r})=\mathbf{z}$, $\delta(\cdot)$ represents the Dirac delta function and $Z=\int\exp{(-\beta U(\mathbf{r}))}\mathrm{d}\mathbf{r}$ is the partition function; we refer to Ref. [41] for details. For high-dimensional CVs, direct estimation of $\rho(\mathbf{z})$ often becomes numerically challenging. An alternative approach is to fit the mean force $\mathbf{F}(\mathbf{z}):=-\nabla A(\mathbf{z})$ at various sample points, which can be estimated via the restrained MD [2] by introducing a harmonic term into the full potential, i.e.,
(3) $U_{k}(\mathbf{r},\mathbf{z})=U(\mathbf{r})+\frac{k}{2}(\mathbf{s}(\mathbf{r})-\mathbf{z})^{\top}(\mathbf{s}(\mathbf{r})-\mathbf{z}),$
@@
where $k$ represents the magnitude of the restrained potential. As shown in Ref [32], the mean force can be computed by $\mathrm{F}(\mathbf{z})=\lim_{k\to\infty}\mathrm{F}^{k}(\mathbf{z})$, where $\mathrm{F}^{k}(\mathbf{z})$ is defined by
(4) $\mathrm{F}^{k}(\mathbf{z})=\frac{1}{Z_{k}(\mathbf{z})}\int k(\mathbf{s}(\mathbf{r})-\mathbf{z})\exp{(-\beta U_{k}(\mathbf{r},\mathbf{z}))}\mathrm{d}\mathbf{r},$
and can be sampled as the first-moment estimation.
@@
In principle, given a collection of sample points, $A(\mathbf{z})$ can be re-constructed (up to a constant) by matching the mean force $-\nabla A(\mathbf{z})$ at the individual points. However, for MD systems with complex energy landscape, the sampling over the phase space could be highly non-trivial due to the prevalence of energy local minima; the training set is often determined a prior as pre-selected points or in a greedy manner. As the number of CVs increases, the empirical random samples often introduce pronounced discretization error. To efficiently construct $A(\mathbf{z})$ in the thermodynamically accessible region, it is desirable to simultaneously optimize the training set and the FES surrogate approximation through certain adaptive sampling based on the posterior residual error. This motivates the present method illustrated below.
2.2 Min-Max formulation
@@
Let $A_{\mathcal{N}}(\mathbf{z})$ denote the NN surrogate of the FES $A(\mathbf{z})$, which is parameterized by minimizing the loss function
(5) $\mathcal{L}_{\mathcal{N}}(\mathbf{z})=\left|\nabla_{\mathbf{z}}A_{\mathcal{N}}(\mathbf{z})+\mathrm{F}(\mathbf{z}))\right|^{2}$
@@
for $\mathbf{z}\in\Gamma$. To solve the minimization problem, we introduce a sampling distribution $q(\mathbf{z})$ and define the weighted loss
(6) $(\mathcal{L}_{\mathcal{N}},q)=\int_{\Gamma}\mathcal{L}_{\mathcal{N}}(\mathbf{z})q(\mathbf{z})\mathrm{d}\mathbf{z}.$
@@
A desired distribution intends to maximize the discrepancy in the dataset for the current NN surrogate $A_{\mathcal{N}}(\mathbf{z})$. Accordingly, we define the maximization problem as
(7) $J[\mathcal{L}_{\mathcal{N}}]=\max_{q}(\mathcal{L}_{\mathcal{N}},q).$
@@
(8) $\min_{A_{\mathcal{N}}}\max_{q}(\mathcal{L}_{\mathcal{N}},q).$
Proposition 2.1.
@@
Assuming that there exists a solution of $\mathcal{L}_{\mathcal{N}}(\mathbf{z})=0$ for $\mathbf{z}\in\Gamma$, then $A_{\mathcal{N}}^{\ast}$ is a solution if and only if it solves (8).
Proof 2.2.
@@
Suppose $A_{\mathcal{N}}^{\ast}$ is the solution for $\mathcal{L}_{\mathcal{N}}(\mathbf{z})=0$, it satisfies $(\mathcal{L}_{\mathcal{N}^{\ast}},q)=0$ for any $q$, i.e., $J[\mathcal{L}_{\mathcal{N}^{\ast}}]=0$. Therefore, $A_{\mathcal{N}}^{\ast}$ is a solution for the minimax problem (8). On the other hand, if $\hat{A}_{\mathcal{N}}$ is the minimizer for problem (8) but not the solution for $\mathcal{L}_{\mathcal{N}}(\mathbf{z})=0$, then there exists $\hat{q}\in V$ such that $(\mathcal{L}_{\hat{\mathcal{N}}},\hat{q})>0$. However, $(\mathcal{L}_{\mathcal{N}^{\ast}},q)=0$ for all $q$, which contradicts the assumption that $\hat{A}_{\mathcal{N}}$ is the minimizer.
@@
Proposition 2.1 shows that direct construction of FES $A(\mathbf{z})$ can be reformulated as an adversarial learning of an optimal solution $A_{\mathcal{N}}(\mathbf{z})$ for the min-max problem (8). Accordingly, the training consists of two components: the minimization step optimizes the NN surrogate with the current training set; the maximization step explores the largest residual region for the current surrogate and essentially establishes an adaptive sampling of the training set based on the posterior residual $\mathcal{L_{N}}(\mathbf{z})$.
@@
To numerically solve the maximization problem, certain regularization needs to be introduced. Otherwise, the maximization problem will simply yield a delta measure, i.e., $\delta(\mathbf{z}-\mathbf{z}^{*})$, where $\mathbf{z}^{\ast}=\arg\max\mathcal{L}_{\mathcal{N}}(\mathbf{z})$. Since the sampling needs to simultaneously identify the max-residual regime and explore the uncharted phase space, we introduce the entropy-based regularization [48, 15] (see also Refs. [19, 34, 43] for gradient-based regularization), and the max-problem is reformulated by
(9) $\min_{q}\int(-\mathcal{L}_{\mathcal{N}}(\mathbf{z})+\kappa_{h}^{-1}\ln q(\mathbf{z}))q(\mathbf{z})\mathrm{d}\mathbf{z}.$
@@
This problem is convex with a unique global minimum at $q^{*}(\mathbf{z})=\exp(-\kappa_{h}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))/Z^{*}$, where $Z^{*}=\int\exp(-\kappa_{h}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))\mathrm{d}\mathbf{z}$ and $\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z})=-\mathcal{L}_{\mathcal{N}}(\mathbf{z})$. The parameter $\kappa_{h}$ is a Lagrangian multiplier, balancing the weight between the max-residual concentration and the scope of exploration, and can be interpreted as the inverse of temperature in statistical physics. An elevated $\kappa_{h}^{-1}$ induces a distribution closer to a uniform distribution. Conversely, a diminished $\kappa_{h}^{-1}$ induces a distribution concentrated near the max-residual point.
Although $q^{*}(\mathbf{z})=\exp(-\kappa_{h}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))/Z^{*}$ solves Eq. (9), we emphasize that this solution is only formal. In particular, the analytical formulation is unknown since it further depends on the residual $\mathcal{L}_{\mathcal{N}}$ (5) and the exact FES $A(\mathbf{z})$. Unfortunately, the numerical evaluation of $\nabla_{\mathbf{z}}A(\mathbf{z}^{\ast})$ at an arbitrary point relies on conducting restrained MD at $\mathbf{z}=\mathbf{z}^{\ast}$, which can be computationally expensive or even thermodynamically inaccessible. This constraint differs from conventional high-dimensional regression or PDE problems, where the residual can be freely queried within the domain. Consequently, common sampling approaches based on the MCMC [37] and Langevin dynamics [38] can not be directly used. Specifically, the effectiveness of the MCMC method relies on the variance of the underlying distribution. However, most configurations in the MD phase space are thermodynamically inaccessible, resulting in large variance and extremely low sampling efficiency. On the other hand, the Langevin-based sampling method depends on $-\nabla_{\mathbf{z}}\ln q^{*}(\mathbf{z})$, which, in turn, relies on computing $\nabla_{\mathbf{z}}\mathcal{L_{N}^{-}}(\mathbf{z})$ and essentially the gradient of the mean force, i.e., $\nabla_{\mathbf{z}}F(\mathbf{z})=-\nabla_{\mathbf{z}}\otimes\nabla_{\mathbf{z}}A(\mathbf{z})$ from restrained MD simulations. However, we can only sample the value of the mean force $F(\mathbf{z})$ rather than its gradient.
Inspired by the consensus-based sampling method [7], we propose a modified interacting particle system governed by the McKean SDE. The gradient-free nature enables us to efficiently construct the mean-field non-local conservative potential and the coherent noise without the evaluation of $\nabla_{\mathbf{z}}F(\mathbf{z})$, and sample the particle distribution as an accurate approximation of $q^{\ast}(\mathbf{z})$ illustrated below.
2.3 Exploitation and exploration in the max-problem